OpenStudy (anonymous):

1. Simplify the following expression. I think the answer is -5y + 5z 3x − 4y + 3z − 3x − y + 2z 6x − 5y + 5z −5y + 5z 6x − 5y + 6z −5y + 6z 2. Which of the following expressions could be simplified further? I think the answer is 6x - 7m + 9x, but I'm not positive. 2x + 3m + 7n + 4z − 5p 6x − 7m + 9x 4x + 3m − 5z + 8a 2x + 3m − 5y + 4z

OpenStudy (anonymous):

What you want to do is to first rearrange the equation so that it has similar letters next to each other: 3x- 4y +3z- 3x- y+ 2z =3x- 3x -4y -y +3z +2z then you just count the result for each letter separately: = (3-3)x + (-4-1)y + (3+2)z =0x + (-5)y +(5)z =-5y+5z

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

I really believe its 2 because you can just add the 6x and the 9x and it wont and the result will be the same.

OpenStudy (anonymous):

any more questions?

OpenStudy (anonymous):

Nope, thanks for the help! I'm no that good at math... :P

OpenStudy (anonymous):

ok. Bye. good luck